July 31, 2017. We are conducting some vibration isolations system test this afternoon and tomorrow morning so the TEM will be unavailable. It should be back online tomorrow afternoon.
Currently browsing category
JEOL JEM-2010 TEM
March 24, 2017. The accelerating voltage system is not working correctly and the TEM cannot be used. I have a call in to the technician. Look for updates as this repair progresses. UPDATE March 27, 2017. The TEM is down bad. There is a failed or partially failed electrical component …
August 31, 2016. The filament in the TEM is unstable. This can be seen as subtle changes in beam intensity and focus. I’ve ordered a new filament to arrive early next week. I’ll replace it on Friday, Sept 9 and 12. The scope can be used in the meantime for …
The TEM went down last night with the power outage. It should be back online this afternoon.
The TEM is back at 200kV. I may set it to 190kV overnight so keep in mind it will take ~1/2 to get it up to 200kV the next day.
The TEM is back on but its not at 200kV yet. I will leave it at 190 over night and run it to 200 tomorrow morning.
All, The TEM is down with water and air problems. I’ll take a look at it this morning and send you an update when I have a better idea of what is wrong. Unless we are lucky I don’t expect it to be back online today.
All of the EDS systems are back online now.
The vacuum leak in the TEM goniometer is repaired. I can begin training new users to insert & remove the holders. Remind me to do that next time you come to use the TEM.
Hitachi service will be here next week to work on the scan problem we have had for several months now. I have blocked out Monday afternoon through noon on Friday. I will let you know if they finish early. JEOL service will be here May 2 to work on the vacuum …